OpenStudy (anonymous):

Hi it's Mom, I need help please. Ready to Give Medal :)The value of -8/15 x 20/64 is: -1/6 1/6 -2/3 2/3 Thanks!! :)

OpenStudy (superdavesuper):

u can break it down to: -8/(3x5) x (4x5)/(8x8)...then u can reduce it to....

OpenStudy (anonymous):

hi. yes i am... i wish i could show him better..

OpenStudy (thebatman):

To be honest, the format this question was given in seems a little confusing. Maybe actually drawing it out, or using the equation function, would assist in interpretation.

OpenStudy (superdavesuper):

-8/15 x 20/64 =-8/(3x5) x (4x5)/(8x4x2) then because 8/8=4/4=5/5=1 =-1/3 x 1/2 =-1/6

OpenStudy (anonymous):

my laptop battery died. thANK u Dave :)

OpenStudy (superdavesuper):

U are welcome Mom :)

OpenStudy (anonymous):

only a few more! lol